QUESTION 3 10 points Save Two blocks of masses mi -2 kg and m2 1.91 kg are connected by a string. One block (m1) slides on a table and the other (m2) hangs from the string, which passes over a pulley. The coefficient of kinetic friction is o.20. What is acceleration of the blocks?
Answer #1

on m1: N - m1 g = 0 (in vertical)

N = m1 g

and f = uk m1 g


inhorizontal, T - f = m1 a

on m2: m2g - T = m2 a

adding,

m2 g - f = (m1 + m2)a


a = (9.8) (1.91 - 0.20x2) / (1.91 + 2)

a = 3.8 m/s^2